the length of a rectangle is 7 inches less than 3 times its width. if the perimeter of the rectangle is 50 inches, what are its dimensions?
Answer by mananth(16949)

width =x
length = 3x-7
..
Perimeter = 50
...
Perimeter = 2*(L+W)
2*(x+3x-7)=50
8x-14=50
8x=50+14
8x=64
x= 8 the width
Length = 3x-7 = 3*8-7
=24-7
=17 length